# PiRD: Physics-informed Residual Diffusion for Flow Field Reconstruction

### v1

@@ 3 Experimental Setup
Experiments were designed to evaluate the effectiveness and robustness of the proposed method, referred to as Physics-informed Residual Diffusion(PiRD), in reconstructing high-fidelity flow fields across various complex and practical scenarios using a two-dimensional Kolmogorov flow dataset. The investigation encompassed multiple tasks under different low-fidelity conditions to simulate potential challenges. The initial set of conditions involved simple low-fidelity scenarios where the high-fidelity flow field, originally at $256\times 256$ pixels, was uniformly down-sampled by a scale of 4 and 8. This represented 4×4$\times$4 × and 8×8$\times$8 × up-sampling tasks, respectively. The second scenario focused on sparse and randomly-distributed data, where only a certain percentage of the data was randomly selected for reconstruction tasks. The sparsity levels considered were $5\%$, and $1.5625\%$ of the original data, aiming to test the capability of reconstructing the high-fidelity flow field from significantly reduced data points. The third scenario combined low-fidelity conditions with varying levels of injected Gaussian noise. For this set of experiments, the low-fidelity flow field was set down-sampled by a scale of 4, and a sparse flow field with $1.5625\%$ of the original data. Gaussian noise was then introduced to these low-fidelity flow fields during the testing stage at densities ranging from $20\%$ to $100\%$. This scenario was designed to assess the resilience of the PiRD method in handling noisy, low-fidelity inputs.
@@ 3.1 Evaluation Metrices
$Loss_{\text{PDE}}(\hat{\omega},\omega)=\frac{1}{n}\sum_{i=1}^{n}|G(\hat{\omega}_{i})|^{2},$ (17)
@@ 3.2 Dataset
where $\omega$ represents vorticity, $u$ is the velocity field, $Re$ is the Reynolds number (set to 1000 for this dataset), and $f(x)$ is a forcing term defined as $f(x)=-4\cos(4x^{2})-0.1\omega(x,t)$, to mitigate energy accumulation at larger scales. The boundary condition of this dataset is periodic, which is also considered in the loss function in our model training. The dataset contains 40 simulations, each with a temporal span of 10 seconds partitioned into 320 frames, and it was down-sampled spatially to a $256\times 256$ grid from a $2048\times 2048$ grid for a more efficient training and less requirement for computational resources.
@@ 3.3 Model Training
The dataset was partitioned into training ($80\%$), validation ($10\%$), and testing ($10\%$) sets. Optimization strategies are based on a cosine annealing learning rate, the maximum number of epochs was set to 200, the maximum and minimum learning rate was set to $1e^{-3}$ and $1e^{-5}$ respectively, $T$ was set to 20, and $\kappa$ was set to 2 to ensure the generalizability of $\omega_{T}$. We set the weights for the data term and PDE term as 0.7 and $2e^{-4}$, respectively. The entire training process was conducted on a single RTX3090 GPU over approximately 3 hours.
@@ 4 Results
4.1 Comparison with current super-resolution methods
In this section, we conduct a quantitative analysis comparing our Physics-informed Residual Diffusion (PiRD) model against two state-of-the-art super-resolution methods: a modified UNet with ResNet block and attention block which is similar to the model proposed by Pant et al.[30], and a physics-informed Denoising Diffusion Probabilistic Model (DDPM) as described by Shu et al. [28]. More specifically, two versions of UNet are trained, one is trained for a direct mapping for a super-resolution task scaling from $64\times 64$ to $256\times 256$, while for the second version of UNet, Gaussian noises are first added to the $64\times 64$ input, then is trained to map to the $256\times 256$ output in order to mimic the process that a denoising diffusion model performs, we therefore denote this UNet as Noisy UNet, both UNet models are trained with the loss function including physical laws. To ensure a fair comparison, our PiRD is also trained for a super-resolution task scaling from $64\times 64$ to $256\times 256$. Notice that, unlike the DDPM-based method—which requires only the high-fidelity vorticity field for training—our PiRD model necessitates prior knowledge of the residual between high-fidelity (HR) and low-fidelity (LR) fields.
@@ 4.1 Comparison with current super-resolution methods
Subsequent tests evaluate the Mean Relative Error (MRE) and Partial Differential Equation (PDE) loss across four different tasks: 4×4$\times$4 × up-sampling, 8×8$\times$8 × up-sampling, $5\%$, and $1.5625\%$ reconstruction to $256\times 256$ resolution. For our training and testing, we first use bicubic interpolation on the up-sampling task and Voronoi tessellation on the reconstruction task to make the input matches the desired dimension of our output ($256\times 256$), we then use nearest interpolation on the up-sampling tasks for testing the result, such that all the testing distributions and patterns are UNSEEN for PiRD, UNet, and Noisy Unet for fair comparison. We also iterate each batch 10 times and take the average of the flow fields as our final prediction. This approach is designed to reduce the influence of any randomness in the predictive outcomes. For training PiRD, $T=20$ steps for sampling in each iteration and $\kappa$ was set to 2. In comparison, results for the physics-informed DDPM are generated using the hyper-parameters recommended by the original authors, with Discrete Denoising Implicit Models (DDIM) sampling steps set at 30 and 60 respectively.
Figure 2: Flow field reconstruction from four different low-fidelity inputs (i.e., 4×4$\times$4 × downsampling, 8×8$\times$8 × downsampling, $5\%$ random sampling, and $1.5625\%$ random sampling). A zoomed-in comparison among PiRD, DDPM-based, and UNet methods for three testing cases.
Table 1: Quantitative Comparison: MRE means the Mean Relative Error, PDE means Partial Differential Equation residual of the predicted vorticity field. The best and second best results are highlighted in bold and underline respectively. Note that the models are trained with 4x up-sampling data only.
@@
PiRD(ours) 0.1178 2.0663 0.3580 2.1594 0.2782 1.6939 0.3857 2.0917
@@
As illustrated in Table 1, UNet model struggles across all the tested scenarios particularly in terms of the PDE loss, reflecting its inability to generalize to unseen low-fidelity datasets encountered during training. Although it achieves the second lowest MRE on the 4x up-sampling task, it performs worst on recovering the underlying physical law. The extremely high PDE loss validates our opinion that the flow field reconstruction cannot be merely seen as a computer vision task, which may result in a successful recovery of element-wise information, but unacceptable physical-wise information. This can be further validated by visualizing the sampled flow fields in Figure 2. As shown, UNet model fails to generate reasonable or smooth flow fields, especially when the low-fidelity input contains randomly distributed measurements (e.g., the $5\%$ and $1.5625\%$ reconstruction tasks).
Conversely, the PiRD, DDPM-based and Noisy UNet methods maintained low and reasonable PDE losses, indicating successful recovery of the underlying physical laws during the reconstruction process while containing a low mean relative error. Notably, PiRD performs best on all scenarios on both MRE and PDE loss, which validates our assumption that all different low-fidelity scenarios can first be mapped into a collective noisy LF distribution and then one can utilize the learned reverse Markov chain to sample the LF flow fields back to the HF distribution.
It is noteworthy that the DDPM-based method’s performance on PDE loss is dependent on the number of sampling steps employed; increasing the steps from 30 to 60 tends to reduce PDE loss but at the expense of higher MRE. This trade-off arises because the DDPM-based method involves subtracting the PDE deviation relative to the vorticity field at each sampling step — a process that is challenging to manage without potentially compromising MRE performance. In contrast, our PiRD model excels in predicting the high-fidelity field with both a remarkably low MRE and an exceptionally low PDE loss, showcasing its effectiveness and robustness in the flow field reconstruction task. For a comprehensive and varied examination of PiRD’s performance, we invite readers to consult the Appendix section, where the reconstructed flow fields with different inputs are presented.
@@ 4.2 Evaluation on kinetic energy spectrum and vorticity distribution
The kinetic energy spectrum is a crucial metric in turbulence research and is used to understand the distribution of energy among different scales within the flow field. It characterizes how kinetic energy varies with the wavenumber, revealing the flow’s energy cascade from larger to smaller scales. It is important to highlight that despite UNet’s mean relative error (MRE) being comparable to that of PiRD for the task of 4x up-sampling, as indicated in Table 1, its performance on the kinetic energy spectrum is relatively poor, as shown in Figure 3(a). Although it performs well in recovering large-scale vortices, it fails to retrieve the vorticity at higher wavenumber. This discrepancy further validates our view that the accurate recovery of element-wise data does not necessarily translate to successfully retrieving physically meaningful information. On the contrary, our method demonstrates a robust capability to recover the energy distribution of the turbulent flow as shown in Figure 3(a). The spectrum obtained by our technique aligns closely with the ground truth across a range of wavenumbers, particularly in the inertial subrange, indicating that our method successfully captures the scale-dependent distribution of kinetic energy. The slight deviation at higher wavenumbers can be attributed to the inherent challenges in resolving the smallest scales of turbulence.
Moreover, the vorticity distribution provides insight into the probability of finding certain vorticity magnitudes within the flow. A method that accurately captures the vorticity distribution from coarse data suggests it can retain critical features of the turbulence structure, such as vortex stretching and tilting mechanisms. The vorticity distribution shown in Figure 3(b) further confirms the efficacy of our proposed approach. The recovered distribution exhibits a strong agreement with the ground truth, especially in capturing the peak and the tails of the distribution, which are indicative of the extreme vorticity events in the flow.
@@ 4.3 Evaluation on noisy low-fidelity vorticity field
In this section, we investigate the robustness of PiRD against a progressive density of Gaussian noise injected into the low-fidelity input, the noise injection and prediction process can be described below
@@
where $\alpha$ ranges from 0.2 to 1.0 by a step size of 0.2, the low-fidelity field is down-sampled from the high-fidelity one by a scale of 4, and randomly selected 1.5625% of the high-fidelity data respectively, and then normalized to Gaussian distribution before noise is injected. Table 2 illustrates the performance of MRE and PDE loss of PiRD and UNet under different noise levels for a 4x and a 1.5625% reconstruction task respectively. As shown in Figure 4, as the injected noise density increases, UNet’s performance degrades significantly. On the contrary, PiRD has stably and accurately predicted the HF field for both scenarios. In addition, as illustrated in Figure 5, the kinetic energy spectrum and vorticity distribution appear identical across all noise densities compared to the case where no noise is injected. This observation suggests that PiRD is highly robust to noise interference. Although there is a slight degradation in MRE as the noise density increases, the performance on PDE loss, vorticity distribution, and kinetic energy spectrum remains consistent.
Table 2: Comparison of UNet and PiRD performance for 4x and $1.5625\%$ reconstruction tasks at various noise levels. The models are trained with clean 4x up-sampling data without any noise.
Noise Level MRE Equation Loss
4x $1.5625\%$ 4x $1.5625\%$
UNet PiRD UNet PiRD UNet PiRD UNet PiRD
0.0 0.1207 0.1178 0.4981 0.3857 259.0664 2.0663 998.8791 2.0917
0.2 0.1560 0.1181 0.5045 0.3881 409.8747 2.0563 1232.7024 2.0917
0.4 0.2247 0.1178 0.5101 0.3893 801.6868 2.1201 1537.9834 2.0837
0.6 0.2990 0.1206 0.5132 0.3901 1350.3967 2.1059 1823.7584 2.0946
0.8 0.3723 0.1245 0.5160 0.3903 2006.1515 2.3082 2093.8465 2.3822
1.0 0.4435 0.1318 0.5190 0.3912 2733.3027 2.7030 2560.6975 2.5433
Figure 4: MRE and PDE loss against different levels of injected noise for PiRD and UNet. The left plot is tested based on the 4x up-sampling task, and the right plot is tested on the 1.5625% reconstruction task.
Figure 5: Comparison of PiRD’s performance on kinetic energy spectrum and vorticity distribution for different densities of noise. The models are trained with clean 4x up-sampling data without any noise.
@@ 5 Conclusion
As the first attempt to apply the ResShift diffusion model for the flow field reconstruction problem, the proposed method still has some limitations. While PiRD demonstrates enhanced accuracy and robustness over CNN-based models when dealing with unseen sparse and noisy measurements, its performance can be further optimized when the training and testing sets are under the same distribution. In the near future, incorporating more controllable physical flow information into diffusion model training, as an alternative to the conventional stacking of Gaussian noises, may be a promising direction for developing a more accurate and generalizable flow generation model.

### v2

@@ 3 Experimental Setup
Figure 2: Downsampling methods for the experiments. (a) illustrates an evenly down-sampling operation. The original data is initially downsampled evenly, followed by nearest interpolation to upsample it to match the desired input dimension of the model. (b) depicts a random selection down-sampling operation. A certain percentage of data points are randomly selected from the original dataset, after which the tesselation method is employed on the sparse dataset to achieve the desired dimension. Gaussian noise is introduced to the downsampled dataset during the noisy flow field reconstruction experiments.
Experiments were designed to evaluate the effectiveness and robustness of the proposed method, referred to as Physics-informed Residual Diffusion(PiRD), in reconstructing high-fidelity flow fields across various complex and practical scenarios using a two-dimensional Kolmogorov flow dataset. The investigation encompassed multiple tasks under different low-fidelity conditions to simulate potential challenges. The initial set of conditions involved simple low-fidelity scenarios where the high-fidelity flow field, originally at $256\times 256$ pixels, was uniformly down-sampled by a scale of 4 and 8. This represented 4×4$\times$4 × and 8×8$\times$8 × up-sampling tasks, respectively. The second scenario focused on sparse and randomly-distributed data, where only a certain percentage of the data was randomly selected for reconstruction tasks. The sparsity levels considered were $5\%$, and $1.5625\%$ of the original data, aiming to test the capability of reconstructing the high-fidelity flow field from significantly reduced data points. The third scenario combined low-fidelity conditions with varying levels of injected Gaussian noise. For this set of experiments, the low-fidelity flow field was set down-sampled by a scale of 4, and a sparse flow field with $5\%$ of the original data. Gaussian noise was then introduced to these low-fidelity flow fields during the testing stage at densities ranging from $20\%$ to $100\%$. This experiment was designed to test the capacity of the PiRD in handling noisy, low-fidelity inputs. The data pre-processing methods described above can be visualized in Figure 2.
@@ 3.1 Evaluation Metrices
$Loss_{\text{PDE}}(\hat{\omega})=\frac{1}{n}\sum_{i=1}^{n}|G(\hat{\omega}_{i})|^{2},$ (15)
@@ 3.2 Dataset
where $\omega$ represents vorticity, $u$ is the velocity field, $Re$ is the Reynolds number (set to 1000 for this dataset), and $f(x)$ is a forcing term defined as $f(x)=-4\cos(4x^{2})-0.1\omega(x,t)$, to mitigate energy accumulation at larger scales. The boundary condition of this dataset is periodic. The dataset contains 40 simulations, each with a temporal span of 10 seconds partitioned into 320 snapshots and a spatial resolution of $256\times 256$ grid.
@@ 3.3 Model Training
The dataset was partitioned into training ($80\%$), validation ($10\%$), and testing ($10\%$) sets. Optimization strategies are based on a cosine annealing learning rate, the maximum number of epochs was set to 20, the maximum and minimum learning rate was set to $1e^{-3}$ and $1e^{-5}$ respectively, $T$ was set to 20, and $\kappa$ was set to 2 to ensure the generalizability of $\omega_{T}$. We set the weights for the data term and PDE term as 0.7 and $2e^{-4}$, respectively. The entire training process was conducted on a single RTX3090 GPU over approximately 3 hours.
@@ 4 Results
4.1 Comparison with current reconstruction methods
In this section, we conduct a quantitative analysis comparing our Physics-informed Residual Diffusion (PiRD) model against three state-of-the-art super-resolution methods: a modified UNet with ResNet block and attention block which is similar to the model proposed by Pant et al.[41], a physics-informed Denoising Diffusion Probabilistic Model (DDPM) proposed by Shu et al. [34], and the Residual Diffusion Model (RDM) proposed by Yue et al. [35].
To ensure a fair comparison, UNet, RDM, and PiRD are all trained for a reconstruction task scaling from $64\times 64$ to $256\times 256$. Notice that, unlike the DDPM-based method—which requires only the high-fidelity vorticity field for training—our PiRD model necessitates prior knowledge of the residual between high-fidelity (HR) and low-fidelity (LR) fields.
@@ 4.1 Comparison with current reconstruction methods
Subsequent tests evaluate the Mean Relative Error (MRE) and Partial Differential Equation (PDE) loss across four different tasks: 4×4$\times$4 × up-sampling, 8×8$\times$8 × up-sampling, $5\%$, and $1.5625\%$ reconstruction to $256\times 256$ resolution. We utilize bicubic interpolation to resize the down-sampled low-fidelity dataset during training, ensuring that the input matches the desired dimensions of our output. Subsequently, for up-sampling tasks, we employ nearest interpolation, and for reconstruction tasks, we utilize Voronoi tessellation to resize the input, maintaining consistency with the desired output dimensions ($256\times 256$). This standardized approach ensures that all testing distributions and patterns remain unseen for UNet, RDM and PiRD, facilitating a fair comparison of their performance. As the reverse process of the diffusion models are initialized with a random sample, we also iterate each batch 10 times and take the average of the flow fields as our final prediction. This approach is designed to reduce the influence of randomness in the predictive outcomes to produce a more stable and reproducible result. By computing the standard deviation over these 10 results, we can also provide the uncertainty of the predicted flow fields. For a visualization of the standard deviation of these 10 iterations and the residual of the prediction and the ground truth, we invite readers to consult the Section C.5 of the Appendix. For training PiRD, $T$ was set to 20 and $\kappa$ was set to 2. For inference, we test PiRD’s performance on both $T=20$ and $T=10$, since our experiment demonstrates that PiRD demonstrates a flexibility on $T$ due to the inherent flexibility of our variance schedule design. Results for the physics-informed DDPM are generated using the hyper-parameters recommended by the original authors, with Discrete Denoising Implicit Models (DDIM) sampling steps set at 30 and 60 respectively. The Residual Diffusion Model (RDM) without the physical loss component is trained for an ablation study, and the hyper-parameters for training the RDM are identical with those for training PiRD.
@@
PiRD - 20(ours) 0.1178 2.0663 0.3616 1.9088 0.2889 1.8443 0.3814 2.1696
PiRD - 10(ours) 0.2033 2.1465 0.4726 2.6937 0.1966 1.9114 0.3733 2.9808
RDM - 20 0.1520 49.8802 0.3617 43.1972 0.3027 39.7595 0.3967 45.2784
@@
Table 1: Quantitative Comparison. MRE means the Mean Relative Error, PDE means Partial Differential Equation residual of the predicted vorticity field. The best and second best results are highlighted in bold and underline respectively. All the models are trained with 4x up-sampling data only. The numbers beside the methods represents the inference step (e.g. “-10” represents the prediction takes 10 steps to sample).
Figure 3: Flow field reconstruction from four different low-fidelity inputs (i.e., 4×4$\times$4 × downsampling, 8×8$\times$8 × downsampling, $5\%$ random selection, and $1.5625\%$ random selection). A zoomed-in comparison among PiRD, DDPM-based, and UNet methods for three testing cases.
As illustrated in Table 1, UNet model struggles across all the tested scenarios particularly in terms of the PDE loss, reflecting its inability to generalize to unseen low-fidelity distributions and patterns which haven’t been encountered during training. Although it achieves the second lowest MRE on the 4x up-sampling task, it performs worst on recovering the underlying physical law. The extremely high PDE loss validates our opinion that the flow field reconstruction cannot be merely seen as a computer vision task, which may result in a successful recovery of element-wise information, but unacceptable physical-wise information. This can be further validated by visualizing the zoomed-in qualitative comparison in Figure 3. As shown, UNet model fails to predict reasonable and smooth flow fields, especially when the low-fidelity input contains randomly distributed measurements (e.g., the $5\%$ and $1.5625\%$ reconstruction tasks).
Conversely, PiRD and DDPM-based method, consistently maintains low and reasonable PDE losses, indicating successful recovery of underlying physical laws during the reconstruction process while also minimizing mean relative errors. Notably, PiRD outperforms other methods across all scenarios in terms of both MRE and PDE loss, confirming our assumption that different low-fidelity scenarios can be collectively mapped into a noisy LF distribution, followed by the use of the learned reverse Markov chain to sample LF flow fields back to the HF distribution. Furthermore, compared to the RDM (Residual Diffusion Model), PiRD demonstrates superior performance not only in terms of PDE loss but also in achieving lower Mean Relative Error (MRE). This outcome highlights the effectiveness of incorporating the physical component into the loss function, ensuring adherence to the underlying governing equation and providing guidance for distribution learning, ultimately resulting in reduced MRE loss.
It is noteworthy that the performance of the DDPM-based method on PDE loss is dependent on the number of sampling steps employed. Specifically, increasing the steps from 30 to 60 tends to reduce PDE loss but at the expense of higher MRE. This trade-off arises because the DDPM-based method involves subtracting the PDE deviation relative to the vorticity field at each sampling step - a process that is challenging to manage without potentially compromising MRE performance.
In contrast, our PiRD model excels in predicting the high-fidelity field with both a remarkably low MRE and an exceptionally low PDE loss, demonstrating its effectiveness and robustness in the flow field reconstruction task. For a comprehensive and varied examination of PiRD’s performance, we invite readers to consult the Section C of the Appendix, where the reconstructed flow fields with different inputs are presented.
@@ 4.2 Evaluation on kinetic energy spectrum and vorticity distribution
The kinetic energy spectrum is a crucial metric in turbulence research and is used to understand the distribution of energy among different scales within the flow field. It characterizes how kinetic energy varies with the wavenumber, revealing the flow’s energy cascade from larger to smaller scales. It is important to highlight that despite UNet’s mean relative error (MRE) being comparable to that of PiRD for the task of 4x up-sampling, as indicated in Table 1, its performance on the kinetic energy spectrum is relatively poor, as shown in Figure 4(a). Although it performs well in recovering large-scale vortices, it fails to retrieve the vorticity at higher wavenumber. This discrepancy further validates our view that the accurate recovery of element-wise data does not necessarily translate to successfully retrieving physically meaningful information. On the contrary, our method demonstrates a robust capability to recover the energy distribution of the turbulent flow as shown in Figure 4(a).
The spectrum obtained by our technique aligns closely with the ground truth across a range of wavenumbers, particularly in the inertial subrange, indicating that our method successfully captures the scale-dependent distribution of kinetic energy. The slight deviation at higher wavenumbers can be attributed to the inherent challenges in resolving the smallest scales of turbulence.
Furthermore, the vorticity distribution offers valuable insight into the likelihood of encountering specific vorticity magnitudes within the flow. An accurate method for capturing this distribution from coarse data suggests its ability to preserve essential characteristics of turbulence structure. These include features such as the distribution’s peak and tails, which provide valuable information about underlying turbulence phenomena.
The obtained distribution as illustrates in Figure 4(b), demonstrates a strong alignment with the ground truth, particularly in accurately representing the distribution’s peak and tails. These aspects are indicative of significant vorticity events within the flow.
@@ 4.3 Evaluation on noisy low-fidelity vorticity field
In this section, we investigate the robustness of PiRD against a progressive density of Gaussian noise injected into the low-fidelity input as shown in Figure 2, the noise injection and prediction process can be described below
@@
where $\alpha$ ranges from 0.2 to 1.0 by a step size of 0.2, the low-fidelity field is down-sampled from the high-fidelity one by a scale of 4, and randomly selected 5% of the high-fidelity data respectively, and then normalized to Gaussian distribution before the Gaussian noise is injected. As shown in Figure 5, as the injected noise density increases, UNet’s performance degrades significantly. On the contrary, PiRD has stably and accurately predicted the HF field for both scenarios. In addition, as illustrated in Figure 6, the kinetic energy spectrum and vorticity distribution appear identical across all noise densities compared to the case where no noise is injected. This observation suggests that PiRD is highly robust to noise interference. Although there is a slight degradation in MRE as the noise density increases, the performance on PDE loss, vorticity distribution, and kinetic energy spectrum remains consistent.
Figure 5: PiRD and UNet’s performances on MRE and PDE under a progressive density of injected Gaussian noise. (a and b) illustrate the comparisons of MRE and PDE loss for PiRD and UNet on the noisy 4x reconstruction task. (c and d) illustrate the comparisons of MRE and PDE loss for PiRD and UNet on the noisy 5% reconstruction task.
Figure 6: Reconstruction performance of PiRD and UNet under a progressive density of injected Gaussian noise. (a and b) illustrate PiRD and UNet ’s performances on Noisy low-fidelity vorticity field reconstruction for 4x and 5% senarios respectively. (c and d) illustrate PiRD’s performance on vorticity distribution and kinetic energy spectrum for 4x reconstruction task. (e and f) illustrate PiRD’s performance on vorticity distribution and kinetic energy spectrum for 5% reconstruction task. Both models are trained with clean 4x up-sampling data without any noise.
@@ 5 Conclusion
As the first attempt to apply the Residual Diffusion Model to the flow field reconstruction problem, the proposed method still exhibits certain limitations. While PiRD demonstrates improved accuracy and robustness compared to CNN-based models when handling unseen sparse and noisy measurements, further optimization of its performance is possible when both the training and testing sets adhere to the exact same distribution. Moreover, a correlation between the standard deviation of multiple model samples and the reconstruction error can be established.
In the near future, exploring the integration of more controllable physical flow information into the diffusion model training process, as an alternative to the conventional addition of Gaussian noises, may represent a promising direction for developing a more accurate and generalizable flow generation model.
@@
Appendix A Detailed mathematical derivation of residual diffusion model
A.1 Derivation of the Forward Process
The conditional probability distribution forward process can be described as follows:
$q(\omega_{t}|\omega_{t-1},\tilde{\omega})$ $=\mathcal{N}(\omega_{t};\omega_{t-1}+\alpha_{t}e,\kappa^{2}\alpha_{t}I),$ $t$ $=1,2,\ldots,T,$ (20)
where its equation form can be therefore described as:
$\omega_{t}$ $=\omega_{t-1}+\alpha_{t}(\omega_{t}-\omega)+\kappa\sqrt{\alpha_{t}}\epsilon_{t}$ (21)
$=\omega_{t-1}+\alpha_{t}e+\kappa\sqrt{\alpha_{t}}\epsilon_{t}$
$=(\omega_{t-2}+\alpha_{t-1}e+\kappa\sqrt{\alpha_{t-1}}\epsilon_{t-1})+\alpha_{t}e+\kappa\sqrt{\alpha_{t}}\epsilon_{t}$
$\ldots$
$=\omega+\sum_{i=1}^{t}\alpha_{i}e+\kappa\sum_{i=1}^{t}\sqrt{\alpha_{i}}\epsilon_{i}$
$=\omega+\sum_{i=1}^{t}(\eta_{i}-\eta_{i-1})e+\kappa\sum_{i=1}^{t}\sqrt{\alpha_{i}}\epsilon_{i},$
$=\omega+\eta_{t}e+\kappa\sum_{i=1}^{t}\sqrt{\alpha_{i}}\epsilon_{i},$
where $\kappa\sum_{i=1}^{t}\sqrt{\alpha_{i}}\epsilon_{i}\sim\mathcal{N}(0,Var(\kappa\sum_{i=1}^{t}\sqrt{\alpha_{i}})I)$, according to the scaling rule of variance, we get get:
$\kappa\sum_{i=1}^{t}\sqrt{\alpha_{i}}\epsilon_{i}$ $\sim\mathcal{N}(0,\kappa^{2}\sum_{i=1}^{t}\alpha_{i}I)$ (22)
$\sim\mathcal{N}(0,\kappa^{2}\eta_{t}I)$
finally, we can express $\omega_{t}$ and $q(\omega_{t}|\omega_{t-1},\tilde{\omega})$ as:
$\omega_{t}$ $=\omega+\eta_{t}e+\kappa\sqrt{\eta_{t}}\epsilon_{i}$ (23)
$q(\omega_{t}|\omega,\tilde{\omega})$ $=\mathcal{N}(\omega_{t};\omega+\eta_{t}e,\kappa^{2}\eta_{t}I),$ $t$ $=1,2,\ldots,T,$
A.2 Derivation of the Backward Process
The entire backward process can be described as:
$p(\omega\mid\tilde{\omega})=\int p(\omega_{T}\mid\tilde{\omega})\prod_{t=1}^{T}p_{\theta}(\omega_{t-1}\mid\omega_{t},\tilde{\omega})\,d\omega_{1:T}$ (24)
where $p(\omega_{T}\mid\tilde{\omega})\propto\mathcal{N}(\omega_{T}\mid\tilde{\omega},\kappa^{2}I)$,The optimization for $\theta$ is achieved by minimizing the negative evidence lower bound:
$\text{min}_{\theta}\sum_{t}D_{KL}[q(\omega_{t-1}\mid\omega_{t},\omega,\tilde{\omega})||p_{\theta}(\omega_{t-1}\mid\omega_{t},\tilde{\omega})]$ (25)
The according to Bayes’s rule, $q(\omega_{t-1}\mid\omega_{t},\omega,\tilde{\omega})\propto q(\omega_{t}\mid\omega_{t-1},\tilde{\omega})q(\omega_{t-1}\mid\omega,\tilde{\omega})$, where
$q(\omega_{t}\mid\omega_{t-1},\tilde{\omega})$ $=\mathcal{N}(\omega_{t};\omega_{t-1}+\alpha_{t}e_{0},\kappa^{2}\alpha_{t}I)$ (26)
$q(\omega_{t-1}\mid\omega,\tilde{\omega})$ $=\mathcal{N}(\omega_{t-1};\omega+\eta_{t-1}e_{0},\kappa^{2}\eta_{t-1}I)$
Then extract the exponent component of the quadratic form of $q(\omega_{t-1}\mid\omega_{t},\omega,\tilde{\omega})$, we can get:
$-\frac{1}{2}\left[\frac{1}{\kappa^{2}\alpha_{t}}+\frac{1}{\kappa^{2}\eta_{t-1}}\right]\omega_{t-1}\omega_{t-1}^{T}+\left[\frac{\omega_{t}-\alpha_{t}e_{0}}{\kappa^{2}\alpha_{t}}+\frac{\omega+\eta_{t-1}e_{0}}{\kappa^{2}\eta_{t-1}}\right]\omega_{t-1}^{T}+C$ (27)
$=-\frac{(\omega_{t-1}-\mu)(\omega_{t-1}-\mu)^{T}}{2\lambda^{2}}+C$
where $\mu=\frac{\eta_{t-1}}{\eta{t}}\omega_{t}+\frac{\alpha_{t}}{\eta_{t}}\omega$ and $\lambda^{2}=\kappa^{2}\frac{\eta_{t-1}}{\eta_{t}}\alpha_{t}$, $C$ is the constant.
A.3 Noise Schedule Design
Recall that $\{\eta_{t}\}_{t=1}^{T}$ is the shifting schedule that increases with the time step $t$ such that a smooth transition of $\omega$ to $\tilde{\omega}$ over a sequence of $T$ steps can be built. The $\eta_{1}$ should satisfies that $\eta_{1}\rightarrow 0$, therefore $\eta_{1}$ is set to be the minimum value between $(0.04/\kappa)^{2}$ and 0.001. The $\eta_{T}$ should satisfies that $\eta_{T}\rightarrow 1$, therefore $\eta_{T}$ is set equal to 0.999. The intermediate time step $\eta_{t}$ is designed as follows:
$\sqrt{\eta_{t}}=\sqrt{\eta_{1}}\times b_{0}^{\beta_{t}},\quad t=2,\dots,T-1$ (28)
where
$\beta_{t}=(\frac{t-1}{T-1})^{p}\times(T-1),b_{0}=exp\left[\frac{1}{2(T-1)}log\frac{\eta_{T}}{\eta_{1}}\right].$ (29)
$p$ is fixed at 0.3 for our experiments.
Appendix B Ablation Studies
B.1 PiRD’s performance under different training configurations
Configurations 4x 8x 5% 1.5625%
T kappa MRE PDE MRE PDE MRE PDE MRE PDE
10 2 0.1198 1.8086 0.3619 1.8025 0.2885 1.6290 0.3856 1.9341
15 2 0.1196 1.9842 0.3601 1.8103 0.2910 1.8470 0.3866 2.0792
20 2 0.1186 2.0727 0.3616 1.9088 0.2889 1.8443 0.3814 2.1696
30 2 0.1224 1.8243 0.3590 1.7667 0.2922 1.7010 0.3866 2.0133
1 0.0847 1.9991 0.3648 1.9357 0.2869 1.8438 0.4045 2.8366
20 3 0.1539 2.1359 0.3627 2.0230 0.2980 2.0774 0.3776 2.1410
4 0.2011 3.3922 0.3795 2.8995 0.3054 3.4197 0.3767 3.2438
Table 2: PiRD’s performance under different training configurations
B.2 PiRD’s performance under different noise level
Noise Level MRE Equation Loss
4x $5\%$ 4x $5\%$
UNet PiRD UNet PiRD UNet PiRD UNet PiRD
0.0 0.1207 0.1178 0.3478 0.2889 259.0664 2.0663 668.0065 1.8443
0.2 0.1560 0.1233 0.3659 0.2906 409.8747 2.0963 851.9433 1.8587
0.4 0.2247 0.1373 0.4022 0.3007 801.6868 2.1556 1332.9699 1.9334
0.6 0.2990 0.1599 0.4533 0.3189 1350.3967 2.2641 1992.5789 2.0817
0.8 0.3723 0.1907 0.5126 0.3458 2006.1515 2.4555 2765.8066 2.3558
1.0 0.4435 0.2302 0.5751 0.3826 2733.3027 2.7867 3618.42871 2.8328
Table 3: Comparison of UNet and PiRD performance for 4x and $5\%$ reconstruction tasks at various noise levels. The models are trained with clean 4x up-sampling data without any noise.
Appendix C Qualitative comparisons
C.1 Qualitative comparisons among methods for 4x upsampling task